OpenStudy (anonymous):

//H̴̙̓͛̊ͮ̒ͣ̽E̠̦̟̝̣̥͂L̿̂͞ͅP̳̊̅ͬ ̄̊̌͌Fͪ̀A̛S̗̒ͮ̑̾T̷̜̺̜̰̻͔̃̃͑̽ͦ̎ͅ show steps if you can!!!// Write the quadratic function c(x)=x^2-8x-17 in vertex form.

OpenStudy (anonymous):

so we must complete the square? do you know how to do that?

OpenStudy (anonymous):

Not really, to be honest...

OpenStudy (anonymous):

\[x^2-x-17=0\] we bring the 17 on the right hand side by adding both sides by 17 \[x^2-8x=17\] Now we divide the coefficient of the x term by 2 and add it to both sides \[x^2-8x+(-\frac{ 8 }{ 2 })^2=17+(-\frac{ 8 }{ 2 })^2\] Now we simplify this into vertex for as follows... \[(x-4)^2=33\] which implies \[y=(x-4)^2-33\] is ur vertex form for that equation

OpenStudy (anonymous):

@arachnidsGrip

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

np :)